OpenStudy (emogirlatemoooocow):

What is the value of the variable that makes the statement true? a + (–7) + 3 = 0 A. –4 B. –3 C. 4 D. 7

OpenStudy (anonymous):

first you should combine the like terms. the -7 and the 3

OpenStudy (anonymous):

what is the equation now?

OpenStudy (emogirlatemoooocow):

10?

OpenStudy (♪chibiterasu):

Nope. You are adding to a negative number.

OpenStudy (♪chibiterasu):

i.e. -2 + 4 = 2 -7 + 4 = -3

OpenStudy (anonymous):

Chibi :L

OpenStudy (♪chibiterasu):

:L

OpenStudy (anonymous):

y you steal my student :(

OpenStudy (♪chibiterasu):

WE ARE WORKING TOGETHER TO FIGHT THE FORCES OF MATH